OpenStudy (anonymous):

Find the equation of the circle with center at (-3, 1) and through the point (2, 13).

OpenStudy (anonymous):

Equation of a circle is (x-h)^2+(y-k)^2=r^2 (h,k)= center r=radius you can plug in the center (h,k) and the point you have (x,y) and solve for the radius then youll have what you need to write the equation

OpenStudy (anonymous):

(2+3)^2+(13-1)^2=r^2

OpenStudy (anonymous):

Solve for r